OpenStudy (anonymous):

A rock is dropped from a cliff. How high is the cliff if the rock falls (1/3) of the total height of the cliff in the last second of it's fall?

OpenStudy (anonymous):

2h=gt^2

OpenStudy (anonymous):

now distance travelled in the last second (t) is h/3

OpenStudy (anonymous):

2h/3= g[(t^2 - (t-1)^2)}

OpenStudy (anonymous):

that should do it

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

I'm not realy sure where you are getting the {t^2-(t-1)^2} to be honest. I understand 2h=gt^2.....but then it looks like you divided that by 3 and then didn't do the same to the otherside ha......wait...maybe I do see what you did..you constrained t so that however long it takes for rock to fall 2/3 the height, it's always 1 less than the total fall time....but i'm still not sure how this will give me the cliff's height ha

OpenStudy (anonymous):

from the first eqn get t, which is is the total time traveled

OpenStudy (anonymous):

then the distance in the last second, is the distance in t seconds - distance in t-1 seconds right?

OpenStudy (anonymous):

so h/3 = [gt^2 - g(t-1)^2] / 2

OpenStudy (anonymous):

so now uve got 2 eqns and two variables t and h

OpenStudy (anonymous):

eliminate t from the system of equations and get h

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

the answer is around 173 meters

OpenStudy (anonymous):

and him, your method is incorrect, btw.

OpenStudy (anonymous):

how com?

OpenStudy (anonymous):

2h/3= g[(t^2 - (t-1)^2)} how did you get this? also, in the last second of its fall, the rock already has an initial velocity. You are discounting that.

OpenStudy (anonymous):

tht gets subtracted neway

OpenStudy (anonymous):

linalg009, is that the correct answer in your book?

OpenStudy (anonymous):

learn to read..i explained it above

OpenStudy (anonymous):

I almost feel like I'd have to trial and error it until I get around the correct time for the last second of the fall. ha....It doesn't have an answer in the book :/

OpenStudy (anonymous):

oops....I mean get the right distance for the last second of the fall ha....i ration of 1/3

OpenStudy (anonymous):

ratio*

OpenStudy (anonymous):

so now the qustn is?

OpenStudy (anonymous):

The question is...can I just plug in different values for the initial height of the drop until I get around 1 sec. for the last third of the fall ha..somehow

OpenStudy (anonymous):

wtdu hav to find out temme clearly

OpenStudy (anonymous):

It's the same question. I'm trying to find out the height of the cliff. I was just trying to think of a different method

OpenStudy (anonymous):

ive told u the method..wat part do u not get? im positive its correct

OpenStudy (anonymous):

What are the two equations? 2h=gt^2 and 2h/3= g[(t^2 - (t-1)^2)}..?

OpenStudy (anonymous):

yeah right

OpenStudy (anonymous):

t=sqrt(2h/g)

OpenStudy (anonymous):

or first u divide the two to make a quadratic eqn in t

OpenStudy (anonymous):

then get the value of t and plug it into the first eqn to get h..got it?

OpenStudy (anonymous):

ya...solving now ha

OpenStudy (anonymous):

yes, him is right. I did not read what he had posted correctly.

OpenStudy (anonymous):

the answer is around 145-150 meters.

OpenStudy (anonymous):

I get 145.52 meters

OpenStudy (anonymous):

yes, thats right.

OpenStudy (anonymous):

Perfect! Thanks for the help him1618 and dhatraditya ha..it makes perfect sense now! ha.